OpenStudy (faolan):

Help? Solve the following equation by completing the square. 6m^2 + 5m -3 = 0

OpenStudy (redneckgirl101):

what do youy think it is

OpenStudy (redneckgirl101):

m=−5 ± 97−−√ 12 m≈0.40407148,−1.23740482

OpenStudy (faolan):

srry had a little meeting. i have no idea how to even do the problem @rEdNeCkGiRl101

OpenStudy (redneckgirl101):

i gave u the answer

RhondaSommer (rhondasommer):

so i would personally change m to x

RhondaSommer (rhondasommer):

\[6x^2+5x-3=0\]

OpenStudy (faolan):

my answer need to be a fraction

OpenStudy (faiqraees):

@Faolan If you want an explanation on its solution let me know.

RhondaSommer (rhondasommer):

yup :)

OpenStudy (faolan):

@FaiqRaees ill take an explanantion pls

RhondaSommer (rhondasommer):

i was trying to explain it to you

OpenStudy (faolan):

oh ok

RhondaSommer (rhondasommer):

^_^ you can choose; them or me; i dont mind

OpenStudy (faiqraees):

okay lets first start by taking 3 to the right hand side

OpenStudy (redneckgirl101):

it is a fraction its m=−5 ± 97−−√ over_12

OpenStudy (redneckgirl101):

over -12 sorry my fault

RhondaSommer (rhondasommer):

Begin completing the square. Divide all terms by 6 the coefficient of the squared term

RhondaSommer (rhondasommer):

\[x^2 + 0.8333333333x - 0.5= 0\]

RhondaSommer (rhondasommer):

move the constant term (0.5) to the right \[ x^2 + 0.83333333x - 0.5 + 0.5 = 0 +0.5\]

RhondaSommer (rhondasommer):

The x term is 0.8333333333x. Take half its coefficient (0.4166666667). Square it (0.1736111111) and add it to both sides.

RhondaSommer (rhondasommer):

\[0.8333333333x + 0.1736111111 + x^2 = 0.5 + 0.1736111111\]

RhondaSommer (rhondasommer):

Combine like terms: \[0.5 + 0.1736111111 = 0.6736111111\]\[x^2 + 0.8333333333x + 0.1736111111= 0.6736111111\]

RhondaSommer (rhondasommer):

Factor a perfect square on the left side: \[(x + 0.4166666667)(x + 0.4166666667) = 0.6736111111\] Calculate the square root of the right side: \[0.82073815\] Break this problem into two subproblems by setting \[(x + 0.4166666667)=0.82073815~and-0.82073815\]

RhondaSommer (rhondasommer):

then you simplify

RhondaSommer (rhondasommer):

:) \[(x+0.4166666667)=0.82073815\]\[(x+0.4166666667)=-0.82073815\]

RhondaSommer (rhondasommer):

id need to see your choices to tell you what was what

RhondaSommer (rhondasommer):

:)

OpenStudy (faolan):

i dont have any answer choices.....its a fill in the blank

RhondaSommer (rhondasommer):

ohh lame

RhondaSommer (rhondasommer):

\[x^2+5/6~x−0.5=0\]

RhondaSommer (rhondasommer):

\[x^2+5/6~x-0.5+0.5=0+0.5\]

RhondaSommer (rhondasommer):

then \[x^2+5/6~x=0.5\]

RhondaSommer (rhondasommer):

\[x^2 + 5/6~x=0.5\] 1/2 of 5/6 = 10/6 or 5/3

RhondaSommer (rhondasommer):

add 5/3 to both sides

OpenStudy (redneckgirl101):

i think he has the answer already

RhondaSommer (rhondasommer):

yes; but he is trying to get it explained

RhondaSommer (rhondasommer):

whatever. ill just go. ^_^ you have your answer anyway right? you dont need my help

OpenStudy (faolan):

il take the explanation. then I might b able to do other problems

OpenStudy (faolan):

orrrrrrrrrrrrrrrrrrrrrrrrrrr

OpenStudy (redneckgirl101):

oh sorry